OpenStudy (anonymous):

Can someone please help me with this. I'm beyond desperate at this point. Use the elimination method to solve the following system of equations.2x - y + z = -32x + 2y + 3z = 23x - 3y - z = -4 (-1, -3, -2) (1, -3, -2) (1, 3, -2) (1, 3, 2)

OpenStudy (anonymous):

only the last equation is equal to a number?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

My apologies for not breaking it down, that's how it comes to me

OpenStudy (anonymous):

So far, yes, I'm writing this down so I can refer back to it

OpenStudy (richyw):

do you know matrix algebra?

OpenStudy (anonymous):

@richyw Not that well, but if I were to look in a book and find out the gist of it, then yes

OpenStudy (richyw):

keep listining to cambell for now, but look it up after!

OpenStudy (anonymous):

@richyw Sure thing :)

OpenStudy (anonymous):

http://i.imgur.com/CMgMz.gif

OpenStudy (anonymous):

she bailed.

OpenStudy (richyw):

wait I didn't read what campbel was saying. This is wrong.

OpenStudy (richyw):

\[2x - y + z = -32x + 2y + 3z = 23x - 3y - z = -4\] gives this system of equations\[2x-y+z=-4\]\[-32x+2y+3z=-4\]\[23x-3y-z=-4\]

OpenStudy (anonymous):

no they are not all equal to -4

OpenStudy (anonymous):

or are they?

OpenStudy (richyw):

the original post shows them all =-4, but it might have been a copy and pasting error. But what is written IS what I just showed.

OpenStudy (anonymous):

does not seem correct

OpenStudy (richyw):

But your initial system is incorrect.

OpenStudy (richyw):

i am assuming it was a typo and will just leave though. but for future warning. What the OP posted implied that they were all equal to -4. Copying and pasting this into wolfram alpha also outputs the answer that I got when I made this assumption!

OpenStudy (anonymous):

let's do algebra homework for everyone who posts a semi-coherent question and then logs off.

OpenStudy (richyw):

yup I see that. I am not saying you don't understand it! I was just saying that the OP posted the incorrect question. The answer to the OP question is given here... (and is not one of the answers) http://www.wolframalpha.com/input/?i=2x+-+y+%2B+z+%3D+-32x+%2B+2y+%2B+3z+%3D+23x+-+3y+-+z+%3D+-4+

OpenStudy (anonymous):

We'll call it Open Study and get the government to pay for it.

OpenStudy (anonymous):

I'm back, my apologies. My Internet went out for a short amount of time. Let me scroll up to see how far we've gotten.

OpenStudy (anonymous):

@Algebraic! I didn't mean to log off, it was my mistake. And that's the way the problem was given to me. No need for bullying.

OpenStudy (anonymous):

It makes sense so far, by the way

OpenStudy (anonymous):

bullying. is that your response to every situation wherein someone on the internet isn't serving your immediate interests?

OpenStudy (anonymous):

@Algebraic! Come again? I have no problem in admitting to my wrongdoing, that is if i am doing wrong. All I did was explain why I suddenly disappeared to the kind people who are helping me solve this problem.

OpenStudy (anonymous):

I'm aware of 'her' past @CliffSedge

OpenStudy (anonymous):

Well, ya' know just so anyone interested can get an insight into what's going on. Enjoy.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

And the problem with needinG feedback in a subject i'm ro

OpenStudy (anonymous):

I'm rocky in is

OpenStudy (anonymous):

@InsanelyChaotic nice name. apropos.

OpenStudy (richyw):

can you post a screen shot of what the question looks like? you are either getting bad questions, or are messing them up when posting them here. The answer to what you posted, is x=8/27 y=104/27 and -20/27. That IS the ONLY correct answer to the question. Math is not based on interpretation. What you have us DOES have those answers. So it is either you or your homework that is messing up somewhere. A screenshot is needed to determine this.

OpenStudy (richyw):

sorry z=-20/27

OpenStudy (anonymous):

@richyw That's part of the problem. That is how I am receiving the questions. The big part of not being able to solve these solutions is not knowing how they are set up from the start.

OpenStudy (anonymous):

@Algebraic! Do you honestly think I would care about what you think of a username? Of all things you could say, you point something as minor as a screen name?

OpenStudy (anonymous):

u mad?

OpenStudy (richyw):

@campbell_st This is what the output displays like on my screen. Does it not look like this for you?

OpenStudy (anonymous):

No, I am not mad at all, I'm almost too calm right now, but thank you.

OpenStudy (anonymous):

post an ss of the page with the problem @InsanelyChaotic

OpenStudy (anonymous):

What is incorrectly written? @campbell_st

OpenStudy (richyw):

I realize that. So we need a screen shot of the original question to determine where the error is!

OpenStudy (anonymous):

He did screen shot it, this is the original one: Use the elimination method to solve the following system of equations.2x - y + z = -32x + 2y + 3z = 23x - 3y - z = -4

OpenStudy (anonymous):

...

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

Okay. Are the all equal to-4? Then the question become a lot easier to solve.

OpenStudy (anonymous):

the questions as it was given to you. not as you posted it to OS.

OpenStudy (anonymous):

I can solve it if you confirm that.

OpenStudy (richyw):

@InsanelyChaotic no what we need is to see a screenshot of what YOU see... What we see gives the answer that I posted.

OpenStudy (anonymous):

@Algebraic! You clearly have nothing to offer here. Can you please just get off of this site since you're not wanting to contribute? I'm not even asking you to at this point. It's difficult to learn a single thing, or grasp any part of this with your sarcastic comments.

OpenStudy (anonymous):

One second, @richyw

OpenStudy (anonymous):

lol no.

OpenStudy (anonymous):

2x−y+z=−32x+2y+3z=23x−3y−z=−4 2x−y+z=−4 −32x+2y+3z=−4 23x−3y−z=−4

OpenStudy (anonymous):

There we go.

OpenStudy (anonymous):

the problem's already been solved for you. Too bad he deleted the soln. Also too bad that many people had to waste their time deciphering your equation spaghetti.

OpenStudy (anonymous):

This is where I'm confused, whether it's equal to -4? That's how I personally see it as.

OpenStudy (anonymous):

post an ss or stop posting.

OpenStudy (anonymous):

If I knew how to decipher it myself, believe me I wouldn't have even thought of going on here. My apologies for the mix up and my appreciation towards all who have spent their time working this with me,

OpenStudy (richyw):

@InsanelyChaotic Copying and pasting is NOT screen capturing. We need to see if the error is occurring BEFORE or AFTER you copy and paste it to open study. I am not trying to be rude, I am trying to help you get quality answers.

OpenStudy (anonymous):

I can't screen capture it; this computer doesn't allow me to

OpenStudy (anonymous):

There is no error. That's how it has been given to me.

OpenStudy (anonymous):

Well, clearly there's an error in deciphering what it's asking, but that's how it is asked. I don't know how to separate the equations properly.

OpenStudy (richyw):

ok on your screen does it look like this \[2x - y + z = -32x + 2y + 3z = 23x - 3y - z = -4 \]or this...\[2x - y + z = -32\;\; x + 2y + 3z = 23\:\:x - 3y - z = -4\]

OpenStudy (richyw):

I mean is there a space between the -32 and the x?

OpenStudy (anonymous):

It looks like the first one, @richyw

OpenStudy (anonymous):

No space

OpenStudy (anonymous):

it's 2x-y+z =-3 2x+2y+3z=2 3x-3y-z=-4

OpenStudy (richyw):

ok then there is no deciphering to be done. These ARE all equal to -4 in the way you are solving the problem. There is obviously a technological issue you are having! You will need to figure these out before we can help you on open study! It's simple as that. If you show your math teacher the issue they will understand why it is a problem.

OpenStudy (anonymous):

I appreciate your efforts nonetheless

OpenStudy (richyw):

no worries, I hope you are able to sort it out soon so you can start doing the math problems!

OpenStudy (anonymous):

I appreciate it very much, really. I didn't mean to waste your time but your efforts were thoughtful.